OpenStudy (anonymous):

how do I work (2.37 x 10^8)^3?

OpenStudy (anonymous):

if memory serves u just add exponents

OpenStudy (anonymous):

You need to multiply the exponents and distribute the outer exponent

OpenStudy (unklerhaukus):

(a.b)^n=a^n.b^n

OpenStudy (anonymous):

(2.37)^3 times 10^24

OpenStudy (anonymous):

oh thank you @RaeDeSol!!!

OpenStudy (anonymous):

You're welcome

OpenStudy (unklerhaukus):

\[(2.37 \times 10^8)^3=(2.37)^3\times(10^8)^3=2.37^3\times10^{8\times3}\]